A plane has a cruising speed of 150 mph when there's no wind at this speed of the plane flew 300 miles with the wind in the same amount of time if you 200 miles against the wind find the speed of the wind.
Answer by jorel555(1290)

Let w represent the speed of the wind. Then:
300/(150+w)=200/(150-w)
45000-300w=30000+200w
15000=500w
w=30 mph as the speed of the wind. ☺☺☺☺